# Forward Reachability for Discrete-Time Nonlinear Stochastic Systems via Mixed-Monotonicity and Stochastic Order

### v1

@@ I Introduction
Forward stochastic reachability for discrete time linear systems has a rich literature which we may divide according to the regularity assumptions used to compute the stochastic reachable set. The most popular collection of assumptions comprises linearity of the system and boundedness of the uncertainty, either as a random variable [2, 3, 4] or as a characteristic function [5, 6]; Other works replace linearity with a weaker assumption of nonlinear structure [7, 8, 9]. Data-driven approaches [10, 11, 12, 13, 14] employ the weakest possible assumptions—typically no more than measurability; however, to ascertain a strong guarantee on the value of the state requires a significant number of samples, without a clear notion whether the data-driven representation is an over or under approximation.
Our interests lie in directly obtaining the forward reach set in a tractable manner in nonlinear systems with order-preserving dynamics. The archetype of such systems is the class of monotone systems [15] prominent in cooperative multi-agent systems, biological systems, and traffic networks [16]. The monotonicity property is particularly useful for computing interval approximations to forward reachable sets, as the tightest interval over-approximation to a reachable set can be computed with two dynamical simulations. Many extensions exist available for systems that do not satisfy the true order-preserving property; there is mixed monotonicity [17, 18, 16] for systems that can be decomposed into increasing and decreasing parts, and sensitivity [19, 20] for extension of mixed-monotonicity into continuous-time. For a general overview of interval reachability, with a particular emphasis on order-preserving properties and their generalizations, see [21, 22, 23].
@@ II-B Mixed Monotone Systems
Lemma 1 (Discrete time forward interval reach sets [25, Theorem 1]).
@@ II-C Stochastic Orders
In order to use mixed-monotonicity methods to solve the stochastic reachability problem posed in Problems 1 and 2, we must lift the notion of mixed monotonicity from the deterministic case to the stochastic case. Our approach, in a nutshell, is to use a notion of stochastic order instead of the componentwise vector partial orders used in the ordinary setting of monotone functions. We briefly introduce the stochastic ordering we use, namely that of [24].
@@
Definition 2 (Upper set [26, Definition II.1.2]).
@@
Definition 3 (Stochastic Order of random vectors [24, Theorem 6.B.1]).
@@
Lemma 2 (Stochastic order through a monotone function [24, Theorem 6.B.16.b]).
@@
This follows similarly to the proof of Lemma 2 [24, Theorem 6.B.16.b]. ∎
@@
Definition 4 (Standard construction of a random vector [24, Sec. 6.B.3]).
@@ III Towards a Stochastic-order Theory of Monotone Dynamics
Having established the existence of stochastic mixed-monotone systems, we turn to the order preservation properties of their trajectories. First, note that the random vectors $\hat{\mathbf{x}}$, $\hat{\mathbf{w}}$, $\mathbf{y}$, and $\mathbf{v}$ and their relationship to $\mathbf{x}$ and $\mathbf{w}$ are absent in Definition 5, which is crucial for an initializing an execution. The following result provides a way to prove that two random vectors are equal.
@@
Follows directly from [24, Theorem 6.B.19]. ∎
@@ V Algorithms for computing forward reachable set of a desired probability via stochastic order
With formalism of stochastic order for mixed-monotone systems and deriving the propagation interval reach sets and the resulting probability of the state residing within it, we arrive at an algorithm which addresses Problem 1.
@@ VII Conclusion and Future Directions
The stochastic extension of the mixed-monotonicity property allows us, as in the deterministic case, to compute interval over-approximations of reachable sets through an efficient simulation-based algorithm. We achieve this result by deriving stochastic order for a mixed monotone stochastic system. We establish the correctness guarantee for the algorithm using our stochastic extension of the mixed monotonicity property, recovering a concentration bound on propagation of a distribution through a monotone mapping. Future work will look to improving the concentration bound in order to mitigate the decrease in confidence over long time horizons. Alternatively, we may modify the algorithm to enlarge the $X_{k}$ sets in order to maintain a constant confidence over time.

### v2

@@ I Introduction
Forward stochastic reachability for discrete time linear systems has a rich literature that we may divide according to the regularity assumptions used to compute the stochastic reachable set. The most popular collection of assumptions comprises linearity of the system and boundedness of the uncertainty, either as a random variable [2, 3, 4] or as a characteristic function [5, 6]; Other works replace linearity with a weaker assumption of nonlinear structure [7, 8, 9, 10, 11].
Data-driven approaches [12, 13, 14, 15, 16] employ the weakest possible assumptions—typically no more than measurability; however, to ascertain a strong guarantee on the value of the state requires a significant number of samples, without a clear notion whether the data-driven representation is an over or under approximation.
Our interests lie in directly obtaining the forward reach set in a tractable manner in nonlinear systems with order-preserving dynamics. The archetype of such systems is the class of monotone systems [17] prominent in cooperative multi-agent systems, biological systems, and traffic networks [18]. The monotonicity property is particularly useful for computing interval approximations to forward reachable sets, as the tightest interval over-approximation to a reachable set can be computed with two dynamical simulations. Many extensions exist available for systems that do not satisfy the true order-preserving property; there is mixed monotonicity [19, 20, 18] for systems that can be decomposed into increasing and decreasing parts, and sensitivity [21, 22] for extension of mixed-monotonicity into continuous-time. The applications we consider in this paper focus on the case of mixed monotonicity. Mixed monotone methods provide mode conservatiev overapproximations than either pure monotone methods or sensitivity methods; but because of the relative restrictiveness of the pure monotonicity assumption, and the difficulties involved in computing the sensitivity functions, mixed monotone methods are the most pragmatic class of methods out of the three. Thus, in this paper we elect to make a trade-off that favors extreme computational efficiency, wide applicability, and ease of construction of algorithms at the cost of introducing conservatism into the reachable set over-approximation. It is beyond the scope of this paper to make a comparison of the potential of all three methods for stochastic generalization: for a general overview of interval reachability, with a particular emphasis on order-preserving properties and their generalizations, see [23, 24, 25].
@@ II-B Mixed Monotone Systems
Lemma 1 (Discrete time forward interval reach sets [27, Theorem 1]).
@@ II-C Stochastic Orders
In order to use mixed-monotonicity methods to solve the stochastic reachability problem posed in Problems 1 and 2, we must lift the notion of mixed monotinicity from the deterministic case to the stochastic case. Our approach, in a nutshell, is to use a notion of stochastic order instead of the componentwise vector partial orders used in the ordinary setting of monotone functions. We briefly introduce the stochastic ordering we use, namely that of [26].
@@
Definition 2 (Upper set [28, Definition II.1.2]).
@@
Definition 3 (Stochastic Order of random vectors [26, Theorem 6.B.1]).
@@
Lemma 2 (Stochastic order through a monotone function [26, Theorem 6.B.16.b]).
@@
This follows similarly to the poof of Lemma 2 [26, Theorem 6.B.16.b]. ∎
@@
Definition 4 (Standard construction of a random vector [26, Sec. 6.B.3]).
@@ III Towards a Stochastic-order Theory of Monotone Dynamics
Having established the existence of stochastic mixed-monotone systems, we turn to the order preservation properties of their trajectories. First, note that the random vectors $\hat{\mathbf{x}}$, $\hat{\mathbf{w}}$, $\mathbf{y}$, and $\mathbf{v}$ and their relationship to $\mathbf{x}$ and $\mathbf{w}$ are absent in Definition 5, which is crucial for an initializing an execution. The following result provides a way prove that two random vectors are equal.
@@
Follows directly from [26, Theorem 6.B.19]. ∎
@@ V Algorithms for computing forward reachable set of a desired probability via stochastic order
With formalism of stochastic order for mixed-monotone systems and deriving the propagation interval reach sets and the resuling probability of the state residing within it, we arrive at an algorithm which addresses Problem 1.
@@ VII Conclusion and Future Directions
The stochastic extension of the mixed-monotonicity property allows us, as in the deterministic case, to compute interval over-approximations of reachable sets through an efficient simulation-based algorithm. We achieve this result by deriving stochastic order for a mixed monotone stochastic system, We establish the correctness guarantee for the algorithm using our stochastic extension of the mixed monotonicity property, recovering a concentration bound on propagation of a distribution through a monotone mapping. Future work will look to improving the concentration bound in order to mitigate the decrease in confidence over long time horizons. Alternatively, we may modify the algorithm to enlarge the $X_{k}$ sets in order to maintain a constant confidence over time.